Using a solubility table, complete and balance the double replacement reaction when FeSO4 (aq) combines with Ba(NO3)2 (aq). Then write a complete ionic equation and a net ionic equation for the reaction.​​​​​
1
Expert's answer
2020-03-04T07:59:16-0500

"FeSO_4(aq) + Ba(NO_3)_2(aq) = BaSO_4(s)+Fe(NO_3)_2(aq)"


"Fe^{2+}(aq)+SO_4^{2-}(aq)+ Ba^{2+}(aq)+2NO_3^-(aq) = BaSO_4(s)+Fe^{2+}(aq)+2NO_3^-(aq)"


"SO_4^{2-}(aq) + Ba^{2+}(aq)= BaSO_4(s)"